0.3(z-10) - 0.5z =-6 I need the process more than anything - nothing seems clear. 1st get rid of parenths, then distrubute?
Answer by stanbon(75887)

0.3(z-10) - 0.5z =-6
On the left side of the equal sign, use the distributive property
to get rid of the parenthesis:
0.3z-0.3*10 - 0.5z = -6
Combine "like" terms on the left side:
-0.2z-3 =-6
Add 3 to both sides to get:
-0.2z = -3
Divide both sides by -0.2 to get:
z=15
----------
Cheers,
Stan H.